OpenStudy (ksmd13):

What is the value of 5 to the power of 4 over 5 to the power of 6? 1 over 25 1 over 5 5 25

OpenStudy (ksmd13):

not sure if it's b or c

OpenStudy (ksmd13):

my answer is 1/5 to the power of 2

OpenStudy (aveline):

\[\frac{ 5^ {4} }{ 5 ^{6} }=\frac{ 1 }{ 5^ {2} }=\frac{ 1 }{ 25 }\]

OpenStudy (ksmd13):

okay thanks! :)